# On the Learn-to-Optimize Capabilities of Transformers in In-Context Sparse Recovery

### v2

@@ 4.3 Transformer Can Provably Perform LISTA-type Algorithms
where $\{\bm{\beta}_{i}^{(1)}\}_{i\in[2N+1]}\in\mathbb{R}^{d}$ are implicit parameter vectors initialized as $\mathbf{0}_{d}$, and $\mathbf{x}_{i}$ is the $i$-th column of the transposed measurement matrix, i.e, $[\mathbf{X}^{\top}]_{:,i}$. We note that a similar embedding structure is adopted in Bai et al., (2024).
Self-attention layer. The self-attention layer takes as input a sequence of embeddings and outputs a sequence of embeddings of the same length. Let the $K$-layer decoder-based Transformer feature four attention heads uniquely indexed as $+1$, $-1$, $+2$, and $-2$. We construct these heads according to the structure specified in Section C.1. This construction ensures that the self-attention layer performs the $\bm{\beta}$ updating inside the soft-thresholding function in Equation 4.1. Furthermore, with our construction, the learnable matrix $\mathbf{D}^{(k)}$ in LISTA-CP becomes context-dependent instead of fixed. In the update rule implemented by the self-attention layer, this matrix becomes $\mathbf{D}^{(k)}=\frac{1}{2N+1}\mathbf{X}(\mathbf{M}^{(k)})^{\top}$, where $\mathbf{M}^{(k)}\in\mathbb{R}^{d\times d}$ is fixed.
@@
where $\mathcal{S}_{\theta^{(k)}}$ is the soft-threshold function. Essentially, the soft-threshold function is effectively implemented by the MLP layer utilizing the $\mathrm{ReLU}$ activation. This can be realized by combining $-\mathrm{ReLU}(x)$, $\mathrm{ReLU}(-x)$, $\mathrm{ReLU}(x-\theta)$, $-\mathrm{ReLU}(-x+\theta)$, and $x$. The implementation details of the MLP layer can be found in Section C.1.
@@
We note that the boundedness assumption over $\mathbf{x}$ and $\bm{\beta}^{*}$ ensures the robustness of the Transformer and prevents it from blowing up under ill conditions. A similar assumption is adopted in Bai et al., (2024). The noiseless assumption is for ease of analysis and is common in the analysis of Transformers  (Ahn et al.,, 2024; Fu et al.,, 2023; Bai et al.,, 2024). We note that the following Theorem 4.1 can be straightforwardly extended to the noisy case when the noise is bounded.
We denote the input sequence to the $k$-th self-attention layer as $\mathbf{H}^{(k)}$, and use $\bm{\beta}_{2n+1}^{(k+1)}$ to represent the vector $[\mathbf{H}^{(k+1)}]_{d+1:2d+1,2n+1}$. Then, we state the following theorem.
@@ 5.1 Sparse Vector Estimation
Similar to the proofs in Chen et al., (2018) and Liu and Chen, (2019), the core step in proving convergence is to ensure that $\mathbf{D}^{(k)}$ exhibits small coherence with $\mathbf{X}$, i.e., $(\mathbf{D}^{(k)})^{\top}\mathbf{X}\approx\mathbf{I}_{d\times d}$. In Chen et al., (2018) and Liu and Chen, (2019), such a $\mathbf{D}^{(k)}$ is obtained by minimizing the generalized mutual coherence to a fixed $\mathbf{X}$. However, this results in poor generalization across different $\mathbf{X}$’s. In our proof, we consider $\mathbf{X}$ as a random matrix and leverage its sub-Gaussian properties to prove that if $\mathbf{D}^{(k)}=\mathbf{X}(\mathbf{M}^{(k)})^{\top}$, where $\mathbf{M}^{(k)}$ is associated with the covariance of $\mathbf{X}$, then $\mathbf{D}^{(k)}$ will have small coherence with $\mathbf{X}$ with high probability. We defer the detailed proof of Theorem 5.1 to Section D.1. ∎
@@
The linear convergence rate demonstrated in Theorem 5.1 is due to the incorporation of curvature information into the update rule. Specifically, the learned matrices $\mathbf{M}^{(k)}$ serve as approximations of the inverse Hessian. By leveraging the statistical properties of the problem, these matrices effectively accelerate convergence, enabling the Transformer to mimic second-order optimization methods. This allows the Transformer to overcome the traditional limitations of first-order methods, which are typically restricted to a sublinear $\mathcal{O}(1/K)$ convergence rate unless additional assumptions are introduced (Nesterov,, 2005; Nemirovskij and Yudin,, 1983).
@@
We note that the Transformer can be slightly modified to exploit certain hidden structures in the in-context sparse recovery tasks. Specifically, if the support of $\bm{\beta}$ lies in a subset $\mathbb{S}\subset[1:d]$ with $S<|\mathbb{S}|\leq d$, then by slightly modifying the parameters of the Transformer to ensure $[\bm{\beta}^{(k)}]_{i}=0$ for all $i\notin\mathbb{S}$, the ICL performance can be improved by replacing all $d$ involved in Theorem 5.1 by $|\mathbb{S}|$. We defer the corresponding result and analysis to Section D.2.
@@ 6 Experimental Results
Problem setup. In all experiments, we adhere to the following steps to generate in-context sparse recovery instances. First, we sample a ground truth sparse vector $\bm{\beta}^{*}$ from a $d=20$ dimensional standard normal distribution, and we fix the sparsity of $\bm{\beta}^{*}$ to be $3$ by randomly setting $17$ entries in $\bm{\beta}^{*}$ to zero. Next, we independently sample $N=10$ vectors form a $d$ dimensional standard normal distribution and then contract the measurement matrix $\mathbf{X}\in\mathbb{R}^{10\times 20}$ (each sampled $d$ dimensional random vector is a row in $\mathbf{X}$). We also sample an additional $\mathbf{x}_{N+1}$ from the $d$-dimensional standard Gaussian distribution. We follow the noiseless setting in Bai et al., (2024) for sparse recovery, i.e., $\mathbf{y}=\mathbf{X}\bm{\beta}^{*}$.
Baselines. The baselines for our experiments include traditional iterative algorithms such as ISTA and FISTA (Beck and Teboulle,, 2009). We also evaluate three classical LISTA-type L2O algorithms: LISTA, LISTA-CP, ALISTA (Gregor and LeCun,, 2010; Chen et al.,, 2018; Liu and Chen,, 2019). For all of these algorithms, we set the number of iterations $K=12$. We generate a single fixed measurement matrix $\mathbf{X}$. For each training epoch, we create $I=50,000$ instances from $50,000$ randomly generated sparse vectors. During inference, we evaluate the pre-trained LISTA-type models under two settings: (1) when the measurement matrix remains identical to that used during pre-training, reported as “Fixed $X$”, and (2) when the measurement matrix is varying through random sampling, reported as “Varying $X$”.
@@
Next, we impose an additional constraint on the support of the sparse vectors. For in-context sparse recovery instances used in both pre-training and testing, we set the support of the sparse vector $\bm{\beta}$ to the first $10$ entries, i.e., $\mathbb{S}=\{1,2,\cdots,10\}$. As observed in Figure 1(b), Small TF and GPT-2 significantly improve their performances in Figure 1(a), while other baseline algorithms do not exhibit significant performance improvements. In Figure 1(b), we present the experimental results for a support-selected version of LISTA-VM, referred to as LISTA-VM-SS. This algorithm is a simple variation of LISTA-VM, where we incorporate prior knowledge of the support by setting all columns in $\mathbf{X}$ whose indices are not in the prior support set to be zero vectors. As we claim in Remark 4 and Corollary D.1, a Transformer could perform this LISTA-VM-SS by utilizing prior knowledge of the support. Our results show that the LISTA-VM-SS achieves an in-context prediction error of approximately $0.07$, which is almost 10×10$\times$10 × better than the standard version of LISTA-VM and is comparable to the prediction error of Small TF. This empirical finding corroborates Remark 4.
@@ Appendix A Additional Related Works
L2O leverages machine learning to develop optimization algorithms, aiming to improve existing methods and innovate new ones. As highlighted by Sucker et al., (2024) and Chen et al., (2022), L2O intersects with meta-learning (also known as “learning-to-learn”) and automated machine learning (AutoML).
Unlike meta-learning, which focuses on enabling models to quickly adapt to new tasks with minimal data by leveraging prior knowledge from diverse tasks (Finn et al.,, 2017; Hospedales et al.,, 2021), L2O aims to improve the optimization process itself by developing adaptive algorithms tailored to specific tasks, leading to faster convergence and enhanced performance in model training (Andrychowicz et al.,, 2016; Li and Malik,, 2016). Thus, while meta-learning enhances task adaptability, L2O refines the efficiency of the optimization process. In contrast, AutoML focuses on model selection, optimization algorithm selection, and hyperparameter tuning (Yao et al.,, 2018); L2O distinguishes itself by its ability to generate new optimization techniques through learned models.
L2O has demonstrated significant potential across various optimization fields and applications. For instance, Andrychowicz et al., (2016) introduced a method where optimization algorithms are learned using recurrent neural networks trained to optimize specific classes of functions. Li and Malik, (2016) proposed learning optimization algorithms through reinforcement learning, utilizing guided policy search to develop optimization strategies. Furthermore, Hruby et al., (2022) applied L2O to address the “minimal problem”, a common challenge in computer vision characterized by the presence of many spurious solutions. They trained a multilayer perceptron model to predict initial problem solutions, significantly reducing computation time.
@@
There exist some works aiming to theoretically understand the ICL mechanism in Transformer through their training dynamics. Ahn et al., (2024); Mahankali et al., (2023); Zhang et al., (2023); Huang et al., (2023) investigate the dynamics of Transformers with a single attention layer and a single head for in-context linear regression tasks. Cui et al., (2024) prove that Transformers with multi-head attention layers outperform those with single-head attention. Cheng et al., (2023) show that local optimal solutions in Transformers can perform gradient descent in-context for non-linear functions. Kim and Suzuki, (2024) study the non-convex mean-field dynamics of Transformers, and Nichani et al., (2024) characterize the convergence rate for the training loss in learning a causal graph. Additionally, Chen et al., 2024a investigate the gradient flow in training multi-head single-layer Transformers for multi-task linear regression. Chen and Li, (2024) propose a supervised training algorithm for multi-head Transformers.
The training dynamics of Transformers for binary classification (Tarzanagh et al., 2023b, ; Tarzanagh et al., 2023a, ; Vasudeva et al.,, 2024; Li et al.,, 2023; Deora et al.,, 2023; Li et al., 2024a, ) and next-token prediction (NTP) (Tian et al., 2023a, ; Tian et al., 2023b, ; Li et al., 2024b, ; Huang et al.,, 2024) have also been studied recently.
@@ Appendix B Table of Notations
$N$ Number of Measurement Vectors, i.e., number of rows of $\mathbf{X}$
@@
$\sigma:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ ReLU Function:$[\sigma(x)]_{i}=x_{i}$ if $x_{i}\geq 0$, $[\sigma(x)]_{i}=0$ if $x_{i}<0$
@@ C.1 Transformer Structure
Consider a model consisting of $K$ Transformer layers, where each layer is equipped with four attention heads. These heads are uniquely indexed as $+1$, $-1$, $+2$, and $-2$ to distinguish their specific roles within the layer.
@@
We prove Proposition 1 is true for all $k\geq 2$ by induction. First, $\mathbf{H}^{(k)}$ for $k=2$ satisfies the condition automatically; therefore, Proposition 1 is true when $k=2$. Then, we demonstrate that if Proposition 1 is true for $k-1$, it remains valid for $k$. For odd values of $i$, the token-wise outputs of the $k$-th attention layer corresponding to the first and second heads satisfy:
@@
Recall that for odd $i$, we assume that Proposition 1 is true for $k-1$, then $\bm{\beta}_{i}^{(k)}$ satisfies
@@
$[\partial\ell_{1}(\bm{\beta})]_{i}=\begin{cases}\{\text{sign}([\bm{\beta}]_{i})\}&\text{if }[\bm{\beta}]_{i}\neq 0,\\ {[-1,1]}&\text{if }[\bm{\beta}]_{i}=0.\end{cases}$
@@
Recall that $10<\gamma^{(k)}(\mathbf{D}_{:,i})^{\top}\mathbf{X}_{:,i}<$ 1, we obtain that
@@
[…]
@@ D.2 Proof of Remark 4
For this layer, we assign one attention head to follow the structure described below. Consider the case where $i=2n-1$ and $n\gg N_{0}$, and denote $[\mathbf{X}^{\top}]_{:,n}$ as $\mathbf{x}_{n}$. For the first attention head, its structure is as follows:
@@
We randomly select a set $\{o_{j}^{\prime}:j\geq 2N_{0}-1\}$ such that |$|\{o_{j}^{\prime}\}|\leq\epsilon$, satisfying the following conditions:
  • If |$|\{o_{j}:o_{j}=1\}|\geq pn-\dfrac{r}{2}$, then
    |$\left|\{o_{j}:o_{j}=1\}\setminus\{o_{j}^{\prime}\}\right|=pn-\dfrac{r}{2};$
  • If |$|\{o_{j}:o_{j}=1\}|<pn-\dfrac{r}{2}$, then
    |$\left|\{o_{j}:o_{j}=1\}\cup\{o_{j}^{\prime}\}\right|=pn-\dfrac{r}{2}.$

### v3

@@ 4.3 Transformer Can Provably Perform LISTA-type Algorithms
where $\{\bm{\beta}_{i}^{(1)}\}_{i\in[2N+1]}\in\mathbb{R}^{d}$ are implicit parameter vectors initialized as $\mathbf{0}_{d}$, and $\mathbf{x}_{i}$ is the $i$-th column of the transposed measurement matrix, i.e, $\mathbf{X}^{\top}]_{:,i}$. We note that a similar embedding structure is adopted in bai2023transformers.
Self-attention layer. The self-attention layer takes as input a sequence of embeddings and outputs a sequence of embeddings of the same length. Let the $K$-layer decoder-based Transformer feature four attention heads uniquely indexed as +1+1, 1-1, +2+2, and 2-2. We construct these heads according to the structure specified in Section C.1. This construction ensures that the self-attention layer performs the $\bm{\beta}$ updating inside the soft-thresholding function in Equation 4.1. Furthermore, with our construction, the learnable matrix $\mathbf{D}^{(k)}$ in LISTA-CP becomes context-dependent instead of fixed. In the update rule implemented by the self-attention layer, this matrix becomes $\mathbf{D}^{(k)}=\frac{1}{2N+1}\mathbf{X}(\mathbf{M}^{(k)})^{\top}$, where $\mathbf{M}^{(k)}\in\mathbb{R}^{d\times d}$ is fixed.
@@
where $\mathcal{S}_{\theta^{(k)}}$ is the soft-threshold function. Essentially, the soft-threshold function is effectively implemented by the MLP layer utilizing the $\mathrm{ReLU}$ activation. This can be realized by combining $\mathrm{ReLU}(x)$, $\mathrm{ReLU}(-x)$, $\mathrm{ReLU}(x-\theta)$, $\mathrm{ReLU}(-x+\theta)$, and $x$. The implementation details of the MLP layer can be found in Section C.1.
@@
We note that the boundedness assumption over $\mathbf{x}$ and $\bm{\beta}^{*}$ ensures the robustness of the Transformer and prevents it from blowing up under ill conditions. A similar assumption is adopted in bai2023transformers. The noiseless assumption is for ease of analysis and is common in the analysis of Transformers  (ahn2023transformers; fu2023transformers; bai2023transformers). We note that the following Theorem 4.1 can be straightforwardly extended to the noisy case when the noise is bounded.
We denote the input sequence to the $k$-th self-attention layer as $\mathbf{H}^{(k)}$, and use $\bm{\beta}_{2n+1}^{(k+1)}$ to represent the vector $\mathbf{H}^{(k+1)}]_{d+1:2d+1,2n+1}$. Then, we state the following theorem.
@@ 5.1 Sparse Vector Estimation
Similar to the proofs in chen2018theoretical and liu2019alista, the core step in proving convergence is to ensure that $\mathbf{D}^{(k)}$ exhibits small coherence with $\mathbf{X}$, i.e., $(\mathbf{D}^{(k)})^{\top}\mathbf{X}\approx\mathbf{I}_{d\times d}$. In chen2018theoretical and liu2019alista, such a $\mathbf{D}^{(k)}$ is obtained by minimizing the generalized mutual coherence to a fixed $\mathbf{X}$. However, this results in poor generalization across different $\mathbf{X}$’s. In our proof, we consider $\mathbf{X}$ as a random matrix and leverage its sub-Gaussian properties to prove that if $\mathbf{D}^{(k)}=\mathbf{X}(\mathbf{M}^{(k)})^{\top}$, where $\mathbf{M}^{(k)}$ is associated with the covariance of $\mathbf{X}$, then $\mathbf{D}^{(k)}$ will have small coherence with $\mathbf{X}$ with high probability. We defer the detailed proof of Theorem 5.1 to Section D.1. ∎
@@
The linear convergence rate demonstrated in Theorem 5.1 is due to the incorporation of curvature information into the update rule. Specifically, the learned matrices $\mathbf{M}^{(k)}$ serve as approximations of the inverse Hessian. By leveraging the statistical properties of the problem, these matrices effectively accelerate convergence, enabling the Transformer to mimic second-order optimization methods. This allows the Transformer to overcome the traditional limitations of first-order methods, which are typically restricted to a sublinear $\mathcal{O}(1/K)$ convergence rate unless additional assumptions are introduced (nesterov2005smooth; nemirovskij1983problem).
@@
We note that the Transformer can be slightly modified to exploit certain hidden structures in the in-context sparse recovery tasks. Specifically, if the support of $\bm{\beta}$ lies in a subset $\mathbb{S}\subset[1:d]$ with $S<|\mathbb{S}|\leq d$, then by slightly modifying the parameters of the Transformer to ensure $\bm{\beta}^{(k)}]_{i}=0$ for all $i\notin\mathbb{S}$, the ICL performance can be improved by replacing all $d$ involved in Theorem 5.1 by $|\mathbb{S}|$. We defer the corresponding result and analysis to Section D.2.
@@ 6 Experimental Results
Problem setup. In all experiments, we adhere to the following steps to generate in-context sparse recovery instances. First, we sample a ground truth sparse vector $\bm{\beta}^{*}$ from a $d=20$ dimensional standard normal distribution, and we fix the sparsity of $\bm{\beta}^{*}$ to be $3$ by randomly setting $17$ entries in $\bm{\beta}^{*}$ to zero. Next, we independently sample $N=10$ vectors form a $d$ dimensional standard normal distribution and then contract the measurement matrix $\mathbf{X}\in\mathbb{R}^{10\times 20}$ (each sampled $d$ dimensional random vector is a row in $\mathbf{X}$). We also sample an additional $\mathbf{x}_{N+1}$ from the $d$-dimensional standard Gaussian distribution. We follow the noiseless setting in bai2023transformers for sparse recovery, i.e., $\mathbf{y}=\mathbf{X}\bm{\beta}^{*}$.
Baselines. The baselines for our experiments include traditional iterative algorithms such as ISTA and FISTA (beck2009fast). We also evaluate three classical LISTA-type L2O algorithms: LISTA, LISTA-CP, ALISTA (gregor2010learning; chen2018theoretical; liu2019alista). For all of these algorithms, we set the number of iterations $K=12$. We generate a single fixed measurement matrix $\mathbf{X}$. For each training epoch, we create $I=50,000$ instances from $50,000$ randomly generated sparse vectors. During inference, we evaluate the pre-trained LISTA-type models under two settings: (1) when the measurement matrix remains identical to that used during pre-training, reported as “Fixed XX”, and (2) when the measurement matrix is varying through random sampling, reported as “Varying XX”.
@@
Next, we impose an additional constraint on the support of the sparse vectors. For in-context sparse recovery instances used in both pre-training and testing, we set the support of the sparse vector $\bm{\beta}$ to the first $10$ entries, i.e., $\mathbb{S}=\{1,2,\cdots,10\}$. As observed in Figure 1(b), Small TF and GPT-2 significantly improve their performances in Figure 1(a), while other baseline algorithms do not exhibit significant performance improvements. In Figure 1(b), we present the experimental results for a support-selected version of LISTA-VM, referred to as LISTA-VM-SS. This algorithm is a simple variation of LISTA-VM, where we incorporate prior knowledge of the support by setting all columns in $\mathbf{X}$ whose indices are not in the prior support set to be zero vectors. As we claim in Remark 4 and Corollary D.1, a Transformer could perform this LISTA-VM-SS by utilizing prior knowledge of the support. Our results show that the LISTA-VM-SS achieves an in-context prediction error of approximately $0.07$, which is almost $10\times$ better than the standard version of LISTA-VM and is comparable to the prediction error of Small TF. This empirical finding corroborates Remark 4.
@@ Appendix A Additional Related Works
L2O leverages machine learning to develop optimization algorithms, aiming to improve existing methods and innovate new ones. As highlighted by sucker2024learningtooptimize and chen2022learning, L2O intersects with meta-learning (also known as “learning-to-learn”) and automated machine learning (AutoML).
Unlike meta-learning, which focuses on enabling models to quickly adapt to new tasks with minimal data by leveraging prior knowledge from diverse tasks (finn2017modelagnostic; hospedales2020metalearning), L2O aims to improve the optimization process itself by developing adaptive algorithms tailored to specific tasks, leading to faster convergence and enhanced performance in model training (andrychowicz2016learning; li2016learning). Thus, while meta-learning enhances task adaptability, L2O refines the efficiency of the optimization process. In contrast, AutoML focuses on model selection, optimization algorithm selection, and hyperparameter tuning (yao2018taking); L2O distinguishes itself by its ability to generate new optimization techniques through learned models.
L2O has demonstrated significant potential across various optimization fields and applications. For instance, andrychowicz2016learning introduced a method where optimization algorithms are learned using recurrent neural networks trained to optimize specific classes of functions. li2016learning proposed learning optimization algorithms through reinforcement learning, utilizing guided policy search to develop optimization strategies. Furthermore, hruby2021learning applied L2O to address the “minimal problem”, a common challenge in computer vision characterized by the presence of many spurious solutions. They trained a multilayer perceptron model to predict initial problem solutions, significantly reducing computation time.
@@
There exist some works aiming to theoretically understand the ICL mechanism in Transformer through their training dynamics. ahn2023transformers; mahankali2023one; zhang2023trained; huang2023incontext investigate the dynamics of Transformers with a single attention layer and a single head for in-context linear regression tasks. cui2024superiority prove that Transformers with multi-head attention layers outperform those with single-head attention. cheng2023transformers show that local optimal solutions in Transformers can perform gradient descent in-context for non-linear functions. kim2024transformers study the non-convex mean-field dynamics of Transformers, and nichani2024transformers characterize the convergence rate for the training loss in learning a causal graph. Additionally, chen2024training investigate the gradient flow in training multi-head single-layer Transformers for multi-task linear regression. chen2024provably propose a supervised training algorithm for multi-head Transformers.
The training dynamics of Transformers for binary classification (tarzanagh2023max; tarzanagh2023transformers; vasudeva2024implicit; li2023theoretical; deora2023optimization; li2024training) and next-token prediction (NTP) (tian2023scan; tian2023joma; li2024mechanics; huang2024non) have also been studied recently.
@@ Appendix B Table of Notations
NN Number of Measurement Vectors, i.e., number of rows of $\mathbf{X}$
@@
[…]
@@ C.1 Transformer Structure
Consider a model consisting of $K$ Transformer layers, where each layer is equipped with four attention heads. These heads are uniquely indexed as +1+1, 1-1, +2+2, and 2-2 to distinguish their specific roles within the layer.
@@
We prove Proposition 1 is true for all $k\geq 2$ by induction. First, $\mathbf{H}^{(k)}$ for $k=2$ satisfies the condition automatically; therefore, Proposition 1 is true when $k=2$. Then, we demonstrate that if Proposition 1 is true for k1k-1, it remains valid for $k$. For odd values of $i$, the token-wise outputs of the $k$-th attention layer corresponding to the first and second heads satisfy:
@@
Recall that for odd $i$, we assume that Proposition 1 is true for k1k-1, then $\bm{\beta}_{i}^{(k)}$ satisfies
@@
[…]
@@
Recall that $0<\gamma^{(k)}(\mathbf{D}_{:,i})^{\top}\mathbf{X}_{:,i}<1$, we obtain that
@@
$A=\Bigg\{$ $\min_{i\in\mathbb{S}}\left|([\mathbf{D}_{n}]_{:,i})^{\top}\mathbf{X}_{1:n,i}\right|\geq\frac{n}{n+2}\left(1-\sqrt{\frac{\log S-\log\delta}{nc}}\right),\quad$
@@ D.2 Proof of Remark 4
For this layer, we assign one attention head to follow the structure described below. Consider the case where i=2n1i=2n-1 and $n\gg N_{0}$, and denote $\mathbf{X}^{\top}]_{:,n}$ as $\mathbf{x}_{n}$. For the first attention head, its structure is as follows:
@@
We randomly select a set $\{o_{j}^{\prime}:j\geq 2N_{0}-1\}$ such that $|\{o_{j}^{\prime}\}|\leq\epsilon$, satisfying the following conditions:
  • If $|\{o_{j}:o_{j}=1\}|\geq pn-\dfrac{r}{2}$, then
    $\left|\{o_{j}:o_{j}=1\}\setminus\{o_{j}^{\prime}\}\right|=pn-\dfrac{r}{2};$
  • If $|\{o_{j}:o_{j}=1\}|<pn-\dfrac{r}{2}$, then
    $\left|\{o_{j}:o_{j}=1\}\cup\{o_{j}^{\prime}\}\right|=pn-\dfrac{r}{2}.$